OpenStudy (lovelyanna):

Find k(8) for the function k(n) = 0.007n. A. 0.0078 B. 0.056 C. 0.56 D. 8.007 Is it B?

OpenStudy (hitaro9):

Yeah, it's B.

OpenStudy (lovelyanna):

Ok, thanks. :) Can you check 2 more?

OpenStudy (hitaro9):

Sure thing

OpenStudy (lovelyanna):

Actually, just one: Find w(7) for the function w(x) = 5x + 2. A.59 B. 45 C. 37 D. 13 C?

OpenStudy (hitaro9):

Yeah. All you do is replace the number with the variable. You've got it right.

OpenStudy (lovelyanna):

Alright, thanks! :)

OpenStudy (hitaro9):

No problem.

OpenStudy (lovelyanna):

Merry x-mas. Have a good day!

OpenStudy (dan815):

CHRISTmas!!

OpenStudy (dan815):

how many times i gotta tell you :)

OpenStudy (raffle_snaffle):

xmas!

OpenStudy (lovelyanna):

What is the value of the expression when a = –6 and b = 5? (a+4b)^2 + 2a^2 A.–1061 B. 124 C. 268 D. –311 Every time I solve this, I get a different answer and none of those answers are from the above. Here Is how I solved it: -6 + (4*5) = 14^2 = 196 2*5 = 10.. 10^2 = 100 196 + 100 = 296 That's what I got.. correct?

OpenStudy (lovelyanna):

@jigglypuff314

OpenStudy (lovelyanna):

Hi Emily! :D Merry x-mas

jigglypuff314 (jigglypuff314):

Hello to you too :) and your mistake was in the 2a^2 not 2a

jigglypuff314 (jigglypuff314):

* not (2a)^2

OpenStudy (lovelyanna):

Oh.. hold on

OpenStudy (lovelyanna):

So a^2 = 5^2 = 25*2 = 50?

jigglypuff314 (jigglypuff314):

a = -6 ?

OpenStudy (lovelyanna):

... my bad.. :O

OpenStudy (lovelyanna):

a^2 = -6^2 = 36*2 = 72?

jigglypuff314 (jigglypuff314):

yep :) then 196 + 72 =

OpenStudy (lovelyanna):

C. 268. Wow.. sorry. I didn't notice :P

jigglypuff314 (jigglypuff314):

you got it :) <3 happy holidays! :P

OpenStudy (lovelyanna):

Thanks! :) And to you too! (btw, I don't celebrate x-mas, I just say it to people because some do :P)

jigglypuff314 (jigglypuff314):

yeah, not everyone is christian :P which is why I say happy holidays :)

OpenStudy (lovelyanna):

Yep! Hmm.. Don't know why I don't say that.. Happy holidays! :D

OpenStudy (lovelyanna):

(6-4)^4 a when a = 5. 80?

OpenStudy (lovelyanna):

@jigglypuff314

jigglypuff314 (jigglypuff314):

yup :)

OpenStudy (lovelyanna):

Ok, thank you! :)

jigglypuff314 (jigglypuff314):

you're welcome :)

OpenStudy (lovelyanna):

I'm gonna go now.. Thanks again! :) (for math and you know what!) Have a good day, bye! ♥

jigglypuff314 (jigglypuff314):

bye :) happy holidays! :) <3

OpenStudy (lovelyanna):

You too! =3 See you soon!

OpenStudy (lovelyanna):

If f(x) = |x + 8|, find f(–11). A. –3 B. 3 C. –19 D. 19 C?

OpenStudy (lovelyanna):

@jigglypuff314

jigglypuff314 (jigglypuff314):

not quite :) | (-11) + 8 | = so first what is -11 + 8 = ?

OpenStudy (lovelyanna):

-3, so its A? I thought it was that at first..

jigglypuff314 (jigglypuff314):

then absolute value the -3

OpenStudy (lovelyanna):

But with | | Should be 3?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (lovelyanna):

Thanks :) I have

OpenStudy (lovelyanna):

oops. I have 4 more*

jigglypuff314 (jigglypuff314):

ok :)

jigglypuff314 (jigglypuff314):

ta-da!!!

OpenStudy (lovelyanna):

If f(x) = |x – 4| + 5, find f(1). A. 10 B. 2 C. 0 D. 8 B?

OpenStudy (lovelyanna):

OMG!! :O That is soooo cute!! =D

OpenStudy (lovelyanna):

Wow.. you are amazing.. :'D

jigglypuff314 (jigglypuff314):

you're welcome :) I emailed the file to you :P f(1) = | (1) – 4| + 5 | -3 | + 5 3 + 5

OpenStudy (lovelyanna):

Oh.. then its 8?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (lovelyanna):

Sorry, I didn't check my email, hold on

OpenStudy (lovelyanna):

Sorry.. I got mixed up.. :/

OpenStudy (lovelyanna):

If f(x) = 9 – 2|x + 2|, find f(–7). A. –1 B. –15 C. 1 D. 19 9 – 2|x + 2| should there be multiplication here? ^ |

jigglypuff314 (jigglypuff314):

yep, that's multiplication there :)

OpenStudy (lovelyanna):

Ok, give me a min :)

OpenStudy (lovelyanna):

Should the equation be like this?: 9-2* | -7+2 | 9-2 = 7 |7 + 2 | = 9

OpenStudy (lovelyanna):

Then 7*9? :/

jigglypuff314 (jigglypuff314):

umm... 9 - 2 | (-7) + 2 | 9 - 2 | - 5 | 9 - 2 * (5)

OpenStudy (lovelyanna):

So.... 9 - 2 = 7 * 5?

jigglypuff314 (jigglypuff314):

not equalling involved? 9 - 2 * 5 9 - 10 -1

OpenStudy (lovelyanna):

Wait a min.. you did 2*5 first? I see....

jigglypuff314 (jigglypuff314):

order of operations PEMDAS

OpenStudy (lovelyanna):

Right, I get it now!

OpenStudy (lovelyanna):

Ugh.. now the graphs -.- I hate graphs! I'm soo bad at them >,<

jigglypuff314 (jigglypuff314):

lol I'll help :)

OpenStudy (lovelyanna):

Thanks! your a life saver! :3

jigglypuff314 (jigglypuff314):

no negatives so it's not upsidedown 2|x + 3| ^ so left 3

OpenStudy (lovelyanna):

D

OpenStudy (lovelyanna):

Oh wait, its not!

OpenStudy (lovelyanna):

A

jigglypuff314 (jigglypuff314):

yep! :)

OpenStudy (lovelyanna):

Yay! =D last one!

OpenStudy (lovelyanna):

Write the equation for the graphed function. http://static.k12.com/calms_media/media/1346000_1346500/1346467/2/670b4d1ed1c240cf5b725290b09d1cd0ea801892/HS_A1G2_S2_116163_Edited.jpg A. f(x) = |x – 3| B. f(x) = |x + 3| C. f(x) = |x| – 3 D. f(x) = |x| + 3

jigglypuff314 (jigglypuff314):

absolute value form: y = m|x - h|+k when vertex is (h, k) so when vertex is (-3, 0) h = -3 y = |x - (-3)| = |x + 3|

OpenStudy (lovelyanna):

So the answer is positive, I knew that! :) B

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (lovelyanna):

Thanks a lot! :) I gtg now

jigglypuff314 (jigglypuff314):

k, see ya later :)

OpenStudy (lovelyanna):

See ya! btw, I'm having soo much fun in Animal jam =D

jigglypuff314 (jigglypuff314):

glad I could help :)

OpenStudy (lovelyanna):

Thanks! :) bye bye!